Kayla wants to make 16 fl. oz. of a 60% acid solution by mixing together a 80% acid solution and pure water. How much of each solution must she use?
Answer by nerdybill(7384)

Kayla wants to make 16 fl. oz. of a 60% acid solution by mixing together a 80% acid solution and pure water. How much of each solution must she use?
.
Let x = amount (oz) of 80% acid
then
16-x = amount (oz) of water
.
.80x = .60(16)
.80x = 9.6
x = 9.6/.80
x = 12 oz of 80% acid solution
.
amount of water:
16-x = 16-12 = 4 oz